Performance Lawn Equipment (PLE) is considering changing their process to make mower blades. They conducted seven production runs and recorded costs. Will the new process save PLE money? Use α = 0.05. (Show all six steps of hypothesis testing.)

Current

New Process

$242.00

$242.00

$176.00

$275.00

$286.00

$199.00

$269.00

$219.00

$327.00

$273.00

$264.00

$265.00

$296.00

$435.00
Answer #1

(1) List all the information

From The Given dta

For Current Year: \bar{x} 1 = 265.714, s1 = 47.801, n1 = 7

For: New Year:     \bar{x} 2 = 272.571, s2 = 77.105, n2 = 7

\alpha = 0.05

(2) State the Null and alternate hypothesis

The Hypothesis:

H0: \mu_1 = \mu_2:

H0: \mu_1 < \mu_2:

This is a Left tailed test

__________________________________

(3) State the test to be used and verify the requirements

We use a 2 sample t test, pooled variances, with df = n1 + n2 - 2 = 7 + 7 - 2 = 12

The requirements are

(1) Samples are simple random samples

(2) The samples are independent of each other

__________________________________________

(4) Compute the standardized test statistic


The Pooled Variance:

S_p^2 = \frac{(n1-1)*s1^2+(n2-1)*s2^2}{n1+n2-2} = \frac{6*47.801^2+6*77.105^2}{7+7-2} = 4115

The Test Statistic:

t = \frac{\bar{x}_1-\bar{x}_2}{\sqrt{S_p^2}(\frac{1}{n1}+\frac{1}{n2})} = \frac{266.714-272.572}{\sqrt{4115*(\frac{1}{7}+\frac{1}{7})}} = -0.17

______________________________________________

(5) Find the p value and compare it to the level of significance.

The p value at t = -0.17) is 0.5665.

p value is > 0.05

_________________________________________________

(6) Make a decision, and then write the final conclusion

Fail To Reject H0. There isn't sufficient evidence at the 95% significance level to conclude that new process will save PLE money.

_____________________________________________________